A parallel resonant circuit has a capacitor 100pF in one branch and the impedance of 100 micro Henry plus a a resistance of 10 ohms in another parallel branch, calculate:

(a) resonant frequency

(b) impendance of the circuit at the resonance and the line current at resonance

(c) Q-quality factor of the circuit


1
Expert's answer
2021-07-23T11:19:02-0400

Resources frequency

"f_r=\frac{1}{2\pi\sqrt{LC}}"

"f_r=\frac{1}{2\pi\sqrt{100\times10^{-12}\times100\times10^{-6}}}"

"f_r=1.6" Hz

Impedance

"Z=\sqrt{R^2+({X_L-X_c})^2}"


"Z=\sqrt{10^2+({2\pi\times1.6\times100\times10^{-6}-\frac{1}{2\pi\times1.6\times100\times10^{-12}}})^2}"

"Z=\sqrt{990466.24+100}=\sqrt{990566.2424}=995\Omega"

C.

"Q=\frac{w_rL}{R}=\frac{10^7\times100\times10^{-6}}{10}=100"